OpenStudy (anonymous):

Given that f(x)=-xh(x) h(-2)=4 h'(-2)=-5 find f'(-2)

OpenStudy (anonymous):

I know I have to do this\[-x \frac{ d }{ dx } h(x)\] but im not getting the awnser

OpenStudy (anonymous):

Here are the choices: a)-6 b)-14 c)-13 d)-8 e)6

jimthompson5910 (jim_thompson5910):

f(x)=-xh(x) d/dx[ f(x) ] = d/dx[ -x*h(x) ] d/dx[ f(x) ] = d/dx[ -x ]*h(x) + (-x)*d/dx[ h(x) ] f ' (x) = -1*h(x) - x*h ' (x) I'll let you finish up

OpenStudy (anonymous):

Thank you so much! I didnt think about using the product rule!

jimthompson5910 (jim_thompson5910):

you're welcome